# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ V.4 Absence of magnetization-induced ingap modes in spin-gapless Mott and Kondo insulators
Figure 13: Electronic excitation from the lowest-energy spin-perturbed states of $S/S_{\rm max}=0.05$ in the 1D HM for $U/t=6$. (a), (b) $A^{\sigma}_{X}(k,\omega)t$ for $|X\rangle$ of the magnetized ground state of $S^{z}=S$ for $\sigma=\downarrow$ [(a)] and $\uparrow$ [(b)]. (c) $A_{X}(k,\omega)t$ for $|X\rangle$ of the lowest-energy state of $S/S_{\rm max}=0.05$ and $S^{z}=0$. Gaussian broadening is used with a standard deviation of $0.1t$. The solid green lines indicate $\omega=0$.
@@ V.5 Spectral function of magnetized states via Shiba transformation
Hence, the spin operators are transformed into the $\eta$ operators:
@@
where $\mu=0$ for ${\cal H}(U)$, $H=0$ for ${\cal H}(-U)$, and the spectral function for ${\cal H}(-U)$ is represented as ${\tilde{A}}^{\lambda}_{X}({\bm{k}},\omega)[=\frac{1}{2}\sum_{\sigma}{\tilde{A}}^{\lambda,\sigma}_{X}({\bm{k}},\omega)]$; ${\tilde{A}}_{X}({\bm{k}},\omega)=\sum_{\lambda}{\tilde{A}}^{\lambda}_{X}({\bm{k}},\omega)$.
@@
The numerical results for $A^{\uparrow}_{X}({\bm{k}},\omega)$ and $A^{\downarrow}_{X}({\bm{k}},\omega)$ for $|X\rangle$ of the state corresponding to the magnetized ground state of $S=S^{z}>0$ for ${\cal H}(U)$ [Figs. 11(a), 11(b), 11(d), 11(e), 11(m), 11(n), 11(p), 11(q), 13(a), and 13(b)] were obtained via Eqs. (110) and (114) using those of ${\tilde{A}}_{X}({\bm{k}},\omega)$ for $|X\rangle$ of the state corresponding to the doped ground state for ${\cal H}(-U)$ in the HMs and PAMs.
@@ V.6 Emergent states corresponding to magnetization-induced states in unmagnetized spin-perturbed systems
The state of $|S=j,S^{z}=0\rangle=({\hat{S}}^{+})^{j}|{\rm GS}\rangle^{S^{z}=-j}=({\hat{S}}^{-})^{j}|{\rm GS}\rangle^{S^{z}=j}$ is the lowest-energy state of $S=j$ for $H=S^{z}=0$, where $|{\rm GS}\rangle^{S^{z}=m}$ denotes the magnetized ground state of $S^{z}=m$, because the magnetized ground states of $S^{z}=\pm j$ are the lowest-energy states of $S=j$ and $S^{z}=\pm j$, provided that phase separation for the magnetization density does not occur at $S^{z}=\pm j$ for $\mu=\delta=0$.
@@ V.7 Remarks on nonzero-$S$ states
In the case of neutron radiation, because a single spin can be excited by a neutron, a macroscopic number $\mathcal{O}(N_{\rm u})$ of neutrons that excite spins (excluding reflected and transmitted neutrons which are irrelevant to the excitations) should be injected every lifetime (relaxation time) of the single-spin excited states to keep the number of excited spins macroscopic, as in a steady state, during the measurement, as mentioned in Sec. IV.6.2. Photons can be used instead of neutrons if they can generate spin excitation spinRIXS .
@@ V.8.1 Selection rules for electronic excitation from a state perturbed by spin fluctuation
Electronic modes can also be induced by spin fluctuation, as in the case of charge fluctuation (Sec. IV.7). As a spin-perturbed state, we consider $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ at half filling at $H=0$ for $\lambda=c$ and $f$. The spin-perturbed state $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of $\bm{q}$}, $N_{\rm h}=0$, and $S^{z}=0$ for the unperturbed ground state $|{\rm GS}\rangle$ with the momentum of $\bm{0}$}, $N_{\rm h}=0$, and $S^{z}=0$. The excitation energy of $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ from $|{\rm GS}\rangle$ is denoted as $e^{{\cal T}^{0}}_{\bm{q}}$.
The state obtained by adding an electron with a momentum $\bm{k}$} to $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=-1$. This state can overlap with $|{\tilde{\cal A}}\rangle_{{\bm{k}}+{\bm{q}}}$ [electron-addition excited states in the conventional bands from $|{\rm GS}\rangle$ (Sec. IV.7.1)] and emerge in the electron-addition spectrum, exhibiting
@@
Similarly, the state obtained by removing an electron with a momentum $\bm{k}$} from $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=1$. This state can overlap with $|{\tilde{\cal R}}\rangle_{-{\bm{k}}+{\bm{q}}}$ [electron-removal excited states in the conventional bands from $|{\rm GS}\rangle$ (Sec. IV.7.1)] and emerge in the electron-removal spectrum, exhibiting
@@ V.8.2 Electronic modes induced by spin fluctuation in spin-gapped Mott and Kondo insulators
Figure 14: Electronic excitation from the spin-perturbed states of small $S/S_{\rm max}$ caused by spin fluctuation in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a), (b)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(d), (e)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(g), (h)], 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(j), (k)], 1D KLM for $J_{\rm K}/t=4$ [(m), (n)], and 2D KLM for $J_{\rm K}/t=8$ [(p), (q)]. (a), (d), (g), (j), (m), (p) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(a)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d)], $A_{X}({\bm{k}},\omega)t$ [(g), (m), (p)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(j)] for $|X\rangle$ of the state corresponding to $\prod_{i}S^{c,z}_{{\bm{q}}_{i}}|{\rm GS}\rangle$ at ${\bm{q}}_{i}=({\bm{0}},\pi)$ in (a) and (d); $\prod_{i}(S^{c,z}_{{\bm{q}}_{i}}-S^{f,z}_{{\bm{q}}_{i}})|{\rm GS}\rangle$ at ${\bm{q}}_{i}={\bm{0}}$ in (g), (j), (m), and (p) for $S/S_{\rm max}=0.1$. Gaussian broadening is used with standard deviations of $0.16t$ [(a), (g)], $0.32t$ [(d), (j)], $0.12t$ [(m)], and $0.24t$ [(p)]. (b), (e), (h), (k), (n), (q) Dispersion relations of electronic excitations from $S^{c,z}_{({\bm{0}},\pi)}|{\rm GS}\rangle$ [(b), (e)] and $(S^{c,z}_{\bm{0}}-S^{f,z}_{\bm{0}})|{\rm GS}\rangle$ [(h), (k), (n), (q)] in the effective theory. The dispersion relations of the emergent modes are $\omega=-\varepsilon^{\cal R}_{-{\bm{k}}_{\parallel}}+e^{{\cal T}^{0}}_{\bm{0}}$ (dotted brown curves), $\omega=\varepsilon^{\cal A}_{{\bm{k}}_{\parallel}}-e^{{\cal T}^{0}}_{\bm{0}}$ (dotted blue curves), $\omega=-\varepsilon^{\bar{\cal R}}_{-{\bm{k}}_{\parallel}}+e^{{\cal T}^{0}}_{\bm{0}}$ (dotted orange curves), and $\omega=\varepsilon^{\bar{\cal A}}_{{\bm{k}}_{\parallel}}-e^{{\cal T}^{0}}_{\bm{0}}$ (dotted purple curves), which are shown in the relevant momentum regimes for comparison with the numerical results. Here, $e^{{\cal T}^{0}}_{\bm{0}}$ denotes the excitation energy of the spin mode at ${\bm{k}}_{\parallel}={\bm{0}}$ [red arrows in (c), (f), (i), (l), (o), and (r)]. The dispersion relations of the other modes and continua in (b), (e), (h), (k), (n), and (q) are the same as those in Figs. 3(g), 3(j), 3(h), 3(k), 3(i), and 3(l), respectively. The up and down arrows on the right sides of (b), (e), (h), (k), (n), and (q) indicate $\omega$-shifts from the unperturbed modes by $e^{{\cal T}^{0}}_{\bm{0}}$ and $e^{{\cal T}^{0}}_{\bm{0}}$, respectively. (c), (f), (i), (l), (o), (r) Dispersion relations of spin excitations in the effective theory. The model parameters in (c), (f), (i), (l), (o), and (r) are the same as those in (b), (e), (h), (k), (n), and (q), respectively. The red arrows indicate the excitation energy of the spin mode at ${\bm{k}}_{\parallel}={\bm{0}}$, i.e., $e^{{\cal T}^{0}}_{\bm{0}}$. The dispersion relations of the modes and continua in (c), (f), (i), (l), (o), and (r) are the same as those in Figs. 4(g), 4(j), 4(h), 4(k), 4(i), and 4(l), respectively. The solid green lines indicate $\omega=0$.
@@ V.8.3 Electronic modes induced by spin fluctuation in the 1D and 2D HMs
Figure 15: Electronic excitation from the spin-perturbed states of small $S/S_{\rm max}$ caused by spin fluctuation in the 1D HM for $U/t=6$ [(a), (b)] and 2D HM for $U/t=9$ [(c)]. (a), (c) $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to $\prod_{i}S^{c,z}_{{\bm{q}}_{i}}|{\rm GS}\rangle$ of $S/S_{\rm max}=0.1$ for ${\bm{q}}_{i}=\frac{\bm{\pi}}{2}$. The magenta and red arrows in (c) indicate $\omega=-e^{{\cal T}^{0}}_{\frac{\bm{\pi}}{2}}+\mu_{+}$ and $\omega=e^{{\cal T}^{0}}_{\frac{\bm{\pi}}{2}}+\mu_{-}$, respectively, where $e^{{\cal T}^{0}}_{\frac{\bm{\pi}}{2}}$ denotes the excitation energy of the spin mode at ${\bm{k}}=\frac{\bm{\pi}}{2}$ in the 2D HM. Gaussian broadening is used with a standard deviation of $0.1t$ [(a), (c)]. (b) Dispersion relations from $S^{c,z}_{\frac{\pi}{2}}|{\rm GS}\rangle$ in the 1D HM. Dotted brown curves indicate $\omega=\varepsilon^{\rm holon}_{k-\frac{\pi}{2}}-e^{{\cal T}^{0}}_{\frac{\pi}{2}}$ (upper curve) and $\omega=-\varepsilon^{\rm holon}_{-k+\frac{\pi}{2}}+e^{{\cal T}^{0}}_{\frac{\pi}{2}}$ (lower curve), where $\varepsilon^{\rm holon}_{k}$ and $e^{{\cal T}^{0}}_{\frac{\pi}{2}}$ denote the excitation energy of the holon mode with a momentum $k$ and that of the spin mode at $k=\frac{\pi}{2}$, respectively. Only the relevant holon modes induced by the spin fluctuation for comparison with the numerical results are shown. The dispersion relations of the other modes and continua in (b) are the same as those in Fig. 1(d). The solid green lines indicate $\omega=0$.
@@ VI.1 Outline
In Sec. VI.5, the relation between doping-induced ingap electronic modes and spin-charge separation of strongly correlated insulators is clarified. Section VI.6 discusses the relation between magnetization-induced ingap electronic modes and spin-charge separation.
@@ VI.2 Conventional band insulators
Figure 16: Electronic, spin, and charge excitation of the noninteracting ($U=0$) ladder HM for $t_{\perp}/t=6$ at $\mu=H=0$. (a) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ for $|X\rangle$ of any eigenstate of the Hamiltonian. The dotted orange lines indicate the bottom of the upper band and top of the lower band; the band gap is the region between the dotted orange lines. (b) $S(k_{\parallel},\pi,\omega)t/3$ and $N(k_{\parallel},\pi,\omega)t/4$. The dotted orange lines indicate the lowest spin- and charge-excitation energies and $\omega=0$; the lowest spin- and charge-excitation energies are the same as the band gap. Gaussian broadening is used with a standard deviation of $0.16t$. The solid green lines indicate $\omega=0$.
@@
for the noninteracting ($U=0$) HM and PAM, respectively, where $\prod_{{\bm{k}}\in{\rm FS}}$ indicates the product over the momentum $\bm{k}$} within the Fermi sea.

### v2

@@ V.4 Absence of magnetization-induced in-gap modes in spin-gapless Mott and Kondo insulators
Figure 13: Electronic excitation from the lowest-energy spin-perturbed states of $S/S_{\rm max}=0.05$ in the 1D HM. (a), (b) $A^{\sigma}_{X}(k,\omega)t$ for $|X\rangle$ of the magnetized ground state of $S^{z}=S$ for $\sigma=\downarrow$ [(a)] and $\uparrow$ [(b)]. (c) $A_{X}(k,\omega)t$ for $|X\rangle$ of the lowest-energy state of $S/S_{\rm max}=0.05$ and $S^{z}=0$. The solid green lines indicate $\omega=0$.
@@ V.5 Spectral function of magnetized states via Shiba transformation
Hence, the spin operators are transformed into the $\eta$ operators,
@@
where $\mu=0$ for ${\cal H}(U)$, $H=0$ for ${\cal H}(-U)$, and the spectral function for ${\cal H}(-U)$ is represented by ${\tilde{A}}^{\lambda}_{X}({\bm{k}},\omega)[=\frac{1}{2}\sum_{\sigma}{\tilde{A}}^{\lambda,\sigma}_{X}({\bm{k}},\omega)]$; ${\tilde{A}}_{X}({\bm{k}},\omega)=\sum_{\lambda}{\tilde{A}}^{\lambda}_{X}({\bm{k}},\omega)$.
@@
The numerical results for $A^{\uparrow}_{X}({\bm{k}},\omega)$ and $A^{\downarrow}_{X}({\bm{k}},\omega)$ for $|X\rangle$ of the state corresponding to the magnetized ground state of $S=S^{z}>0$ for ${\cal H}(U)$ [Figs. 11(a), 11(b), 11(d), 11(e), 11(m), 11(n), 11(p), 11(q), 13(a), and 13(b)] were obtained via Eqs. (112) and (116) using those of ${\tilde{A}}_{X}({\bm{k}},\omega)$ for $|X\rangle$ of the state corresponding to the doped ground state for ${\cal H}(-U)$ in the HM and PAM.
@@ V.6 Emergent states corresponding to magnetization-induced states in unmagnetized spin-perturbed systems
The state of $|S=j,S^{z}=0\rangle=({\hat{S}}^{+})^{j}|{\rm GS}\rangle^{S^{z}=-j}=({\hat{S}}^{-})^{j}|{\rm GS}\rangle^{S^{z}=j}$ is the lowest-energy state of $S=j$ for $H=S^{z}=0$, where $|{\rm GS}\rangle^{S^{z}=m}$ denotes the magnetized ground state of $S^{z}=m$, because the magnetized ground states of $S^{z}=\pm j$ are the lowest-energy states of $S=j$ and $S^{z}=\pm j$, provided that phase separation in the magnetization density does not occur at $S^{z}=\pm j$ for $\mu=\delta=0$.
@@ V.7 Remarks on nonzero-$S$ states
In the case of neutron radiation, because a single spin can be excited by a neutron, a macroscopic number $\mathcal{O}(N_{\rm u})$ of neutrons that excite spins (excluding reflected and transmitted neutrons that are irrelevant to the excitations) should be injected every lifetime (relaxation time) of the single-spin excited states to keep the number of excited spins macroscopic, as in a steady state, during the measurement, as mentioned in Sec. IV.6.2. Photons can be used instead of neutrons if they can generate spin excitation spinRIXS .
@@ V.8.1 Selection rules for electronic excitation from a state perturbed by spin fluctuation
Electronic modes can also be induced by spin fluctuation, as in the case of charge fluctuation (Sec. IV.7). As a spin-perturbed state, we consider $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ at half filling at $H=0$ for $\lambda=c$ and $f$. The spin-perturbed state $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of $\bm{q}$}, $N_{\rm h}=0$, and $S^{z}=0$ for the unperturbed ground state $|{\rm GS}\rangle$ with the momentum of $\bm{0}$}, $N_{\rm h}=0$, and $S^{z}=0$. The excitation energy of each eigenstate in $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ from $|{\rm GS}\rangle$ is denoted by $e^{{\cal T}^{0}}_{\bm{q}}$.
The state obtained by adding an electron with a momentum $\bm{k}$} to $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=-1$. Since this state can overlap with $|{\tilde{\cal A}}\rangle_{{\bm{k}}+{\bm{q}}}$ [electron-addition excited states in the conventional bands from $|{\rm GS}\rangle$ (Sec. IV.7.1)], electronic states can emerge in the electron-addition spectrum, exhibiting
@@
Similarly, the state obtained by removing an electron with a momentum $\bm{k}$} from $S^{\lambda,z}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=1$. Since this state can overlap with $|{\tilde{\cal R}}\rangle_{-{\bm{k}}+{\bm{q}}}$ [electron-removal excited states in the conventional bands from $|{\rm GS}\rangle$ (Sec. IV.7.1)], electronic states can emerge in the electron-removal spectrum, exhibiting
@@ V.8.2 Electronic modes induced by spin fluctuation in spin-gapped Mott and Kondo insulators
Figure 14: Electronic excitation from the spin-perturbed states of small $S/S_{\rm max}$ caused by spin fluctuation in the ladder HM [(a), (b)], bilayer HM [(d), (e)], 1D PAM [(g), (h)], 2D PAM [(j), (k)], 1D KLM [(m), (n)], and 2D KLM [(p), (q)]. (a), (d), (g), (j), (m), (p) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(a)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d)], $A_{X}({\bm{k}},\omega)t$ [(g), (m), (p)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(j)] for $|X\rangle$ of the state corresponding to $\prod_{i}S^{c,z}_{{\bm{q}}_{i}}|{\rm GS}\rangle$ at ${\bm{q}}_{i}=({\bm{0}},\pi)$ in (a) and (d); $\prod_{i}(S^{c,z}_{{\bm{q}}_{i}}-S^{f,z}_{{\bm{q}}_{i}})|{\rm GS}\rangle$ at ${\bm{q}}_{i}={\bm{0}}$ in (g), (j), (m), and (p) for $S/S_{\rm max}=0.1$. (b), (e), (h), (k), (n), (q) Dispersion relations of electronic excitations from $S^{c,z}_{({\bm{0}},\pi)}|{\rm GS}\rangle$ [(b), (e)] and $(S^{c,z}_{\bm{0}}-S^{f,z}_{\bm{0}})|{\rm GS}\rangle$ [(h), (k), (n), (q)] in the effective theory. The dispersion relations of the emergent modes are $\omega=-\varepsilon^{\cal R}_{-{\bm{k}}_{\parallel}}+e^{{\cal T}^{0}}_{\bm{0}}$ (dotted brown curves), $\omega=\varepsilon^{\cal A}_{{\bm{k}}_{\parallel}}-e^{{\cal T}^{0}}_{\bm{0}}$ (dotted blue curves), $\omega=-\varepsilon^{\bar{\cal R}}_{-{\bm{k}}_{\parallel}}+e^{{\cal T}^{0}}_{\bm{0}}$ (dotted orange curves), and $\omega=\varepsilon^{\bar{\cal A}}_{{\bm{k}}_{\parallel}}-e^{{\cal T}^{0}}_{\bm{0}}$ (dotted purple curves), which are shown in the relevant momentum regimes for comparison with the numerical results. Here, $e^{{\cal T}^{0}}_{\bm{0}}$ denotes the excitation energy of the spin mode at ${\bm{k}}_{\parallel}={\bm{0}}$ [red arrows in (c), (f), (i), (l), (o), and (r)]. The dispersion relations of the other modes and continua in (b), (e), (h), (k), (n), and (q) are the same as those in Figs. 3(g), 3(j), 3(h), 3(k), 3(i), and 3(l), respectively. The up and down arrows on the right sides of (b), (e), (h), (k), (n), and (q) indicate $\omega$-shifts from the unperturbed modes by $e^{{\cal T}^{0}}_{\bm{0}}$ and $e^{{\cal T}^{0}}_{\bm{0}}$, respectively. (c), (f), (i), (l), (o), (r) Dispersion relations of spin excitations in the effective theory in the ladder HM [(c)], bilayer HM [(f)], 1D PAM [(i)], 2D PAM [(l)], 1D KLM [(o)], and 2D KLM [(r)]. The red arrows indicate the excitation energy of the spin mode at ${\bm{k}}_{\parallel}={\bm{0}}$, i.e., $e^{{\cal T}^{0}}_{\bm{0}}$. The dispersion relations of the modes and continua in (c), (f), (i), (l), (o), and (r) are the same as those in Figs. 4(g), 4(j), 4(h), 4(k), 4(i), and 4(l), respectively. The solid green lines indicate $\omega=0$.
@@ V.8.3 Electronic modes induced by spin fluctuation in the 1D and 2D HMs
Figure 15: Electronic excitation from the spin-perturbed states of small $S/S_{\rm max}$ caused by spin fluctuation in the 1D HM [(a), (b)] and 2D HM [(c)]. (a), (c) $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to $\prod_{i}S^{c,z}_{{\bm{q}}_{i}}|{\rm GS}\rangle$ of $S/S_{\rm max}=0.1$ for ${\bm{q}}_{i}=\frac{\bm{\pi}}{2}$. The magenta and red arrows in (c) indicate $\omega=-e^{{\cal T}^{0}}_{\frac{\bm{\pi}}{2}}+\mu_{+}$ and $\omega=e^{{\cal T}^{0}}_{\frac{\bm{\pi}}{2}}+\mu_{-}$, respectively, where $e^{{\cal T}^{0}}_{\frac{\bm{\pi}}{2}}$ denotes the excitation energy of the spin mode at ${\bm{k}}=\frac{\bm{\pi}}{2}$ in the 2D HM. (b) Dispersion relations from $S^{c,z}_{\frac{\pi}{2}}|{\rm GS}\rangle$ in the 1D HM. Dotted brown curves indicate $\omega=\varepsilon^{\rm holon}_{k-\frac{\pi}{2}}-e^{{\cal T}^{0}}_{\frac{\pi}{2}}$ (upper curve) and $\omega=-\varepsilon^{\rm holon}_{-k+\frac{\pi}{2}}+e^{{\cal T}^{0}}_{\frac{\pi}{2}}$ (lower curve), where $\varepsilon^{\rm holon}_{k}$ and $e^{{\cal T}^{0}}_{\frac{\pi}{2}}$ denote the excitation energy of the holon mode with a momentum $k$ and that of the spin mode at $k=\frac{\pi}{2}$, respectively. Only the relevant holon modes induced by the spin fluctuation for comparison with the numerical results are shown. The dispersion relations of the other modes and continua in (b) are the same as those in Fig. 1(d). The solid green lines indicate $\omega=0$.
@@ VI.1 Outline
In Sec. VI.5, the relation between doping-induced in-gap electronic modes and spin-charge separation of strongly correlated insulators is clarified. Section VI.6 discusses the relation between magnetization-induced in-gap electronic modes and spin-charge separation.
@@ VI.2 Conventional band insulators
Figure 16: Electronic, spin, and charge excitations of the noninteracting ($U=0$) ladder HM for $t_{\perp}/t=6$ at $\mu=H=0$. (a) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ for $|X\rangle$ of any eigenstate of the Hamiltonian. The dotted orange lines indicate the bottom of the upper band and top of the lower band; the band gap is the region between the dotted orange lines. (b) $S(k_{\parallel},\pi,\omega)t/3$ and $N(k_{\parallel},\pi,\omega)t/4$. The dotted orange lines indicate the lowest spin- and charge-excitation energies and $\omega=0$; the lowest spin- and charge-excitation energies are the same as the band gap. The solid green lines indicate $\omega=0$.
@@
for the noninteracting ($U=0$) HM and PAM, respectively, where $\prod_{{\bm{k}}\in{\rm FS}}$ indicates the product over the momenta $\bm{k}$} within the Fermi sea.